OpenStudy (anonymous):

fifteen years ago a national park had a fox population of 325. Today the fox population is 650 assuming that the population experienced exponential growth, write a relation repersenting the population size of the foxes in the park.this is part b use the relation from the previous question to project the fox population in 20 years.

OpenStudy (anonymous):

doubles every 15 years so using t as time in years you can write \[325\times 2^{\frac{t}{15}}\] where t starts 15 years ago

OpenStudy (anonymous):

or if you prefer \[650\times 2^{\frac{t}{15}}\] if time starts now

OpenStudy (anonymous):

therefore in 20 years from now it will be about \[650\times 2^{\frac{20}{15}}=650\times 2^{\frac{4}{3}}\] then a calculator